The integrating factor of the given DE are already below. What is the value of "n".


(2y^2)dx-x((2x^3)+y)dy=0


IF=y^n


1
Expert's answer
2021-06-25T07:06:04-0400

Exact differential equation,

"y^n[(2y^2)dx-x((2x^3)+y)dy]=0\\
Where,\\
M(x,y)=2y^{2+n}\\
N(x,y)=-2x^4y^n-xy^{1+n}\\
Then,\\
M_y=N_x\\
2(2+n)y^{1+n}=-8x^3y^n-y^{1+n}\\
(4+2n)y^{1+n}=-8x^3y^n-y^{1+n}\\
\text{There doesn't exist any n for which}y^n\text{is a integrating factor.}"